Assuming that there are no heat effects and frictional effects. find the kinetic energy and speed of a 3220 lb body after it falls 778 ft from rest.


1
Expert's answer
2021-10-24T18:29:59-0400

Assuming that there are no heat effects and no factional effects, find the kinetic energy and speed of a 3220-lb body after it falls 778 ft from rest. Start with the steady flow equation, deleting energy terms which are irrelevant.

"m=3220 \;lb \\

h=778 \;ft \\

u=0 \\

v^2=0^2+2gh \\

v^2 = 2gh \\

v = \sqrt{2gh} \\

g=9.8 \;m/s^2 = 9.8 \times 3.28 = 32.14 \;ft/s^2 \\

v = \sqrt{2 \times 32.14 \times 778} \\

v= 223.64 \;ft/s \\

v≈ 224 \; ft/s \\

Kinetic \; energy \;KE = \frac{1}{2}mv^2 \\

KE = \frac{3200 \times (224 )^2}{2 \times 32.2 } \\

KE = 2510827.376 \;lb \; ft \\

KE = \frac{2510827.376}{778} = 3227.285 \;BTU"